OpenStudy (anonymous):

I need to rationalize the denominator and simplify this equation: 1 - the square root of 5 over 6 + 2 times the square root of 5. The first thing I did was multiply everything by 6 - 2 times the square root of 5 and now I'm stuck. Please help me

OpenStudy (anonymous):

\[\large =\frac {1- \sqrt {5}}{6+2 \sqrt {5} }\] that the problem?

OpenStudy (anonymous):

don't mind the = sign

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what u did is correct...

OpenStudy (anonymous):

 can you take it from there?